When interpreting F (2, 27) = 8.80, p < .05, how many total participants were there in the study?

26

29

28

27

30
Answer #1

Solution:

From given information, we have

Total df = 2 + 27 = 29

We have

Total number of participants = total df + 1 = 29 + 1 = 30

Answer: 30